For the year, Beta had beginning total liabilities of $25,000 and ending total liabilities of S15,000. During the year, total assets increased s30.000. The begimnmg balance in stockholders equity was $40,000. In addition, common stock of $10,000 was sold at par value, and dividends of $5,000 were declared and paid. How much was Betas net income for the year. HINT: Use changes in the accounting equation to solve this problem. A. S5,000 B. $10,000 C. $15,000 D. $20,000 E. None of the above If the balance in the accounts payable account decreased during the year, which of the following best explains the adjustment made when converting from cash basis net income to accrual basis net income? A. Because this decrease is the result of recognizing an expense for accrual basis but not cash basis and, therefore, this decrease is required to be added to the cash basis net income balance to find accrual basis net income. B. Because this decrease is the result of recognizing an expense for cash basis but not accrual basis and, therefore, this decrease is required to be subtracted from the cash basis net imcome balance to find accrual basis net income. C. Because this decrease is the result of recognizing an expense for accrual basis but not cash basis and, therefore, this decrease is required to be subtracted from the cash basis net imcome balance to fmd accrual basis net income. D. Because this decrease is the result of recognizing an expense for cash basis but not accrual basis and, therefore, this decrease is required to be added to the cash basis net income balance to find accrual basis net income. E. None of the above Alpha had net sales of S16 million, ending inventory of S6 million, beginnmg inventory of $8 million, and a gross profit percent of 25%. Using the periodic inventory system, what was Apha et purchases? A. $6 million B. S8 million C. $10 million D. $12 million E. None of the above Alpha pays rent on numerous offices located throughout the city each year. Alpha pays rents in advance in some cases, and im others, rents are paid when due. The following information is imcluded in Alphas December 31 balance sheets for 2017 and 2018 017 EI0 Rents Rent Payable $25.000 20,000 During 2018, Alpha paid total cash of S80,000 in rentals. In its income statement for the year ended December 31, 2018, hovw much rent expense should Alpha report on an accrual basis? A. $55,000 B. S65,000 C. $95,000 D. S105,000 E. None of the above Alpha recorded the purchase of supplies as a debit to Supplies Expense for $1,200 and a credit to Cash for S1.200. During the year, only S900 of the supplies were used. This was the only purchase of supplies during the year, and the supplies account had a zero beginnimg balance. The adjusting journal entry that should be recorded at the end of the year is A. Debit Supplies Expense for $300 and credit Supplies for $300 B. Debit Supplies for S300 and credit Supplies Expense for $300. C. Debit Supplies for $900 and credit Supplies Expense for $900 D. Debit Supplies Expense for S900 and credit Supplies for $900. E. None of the above
Answer #1

1)Assets=liabilites+equity
Assets incereased by 30000 so liab+equity also should increase by 30000
liabilites decreased by 10000(25000-15000) so equity has to increase by 30000+10000=40000
Beg equity=40000
ending equity=40000+40000=80000
Net income=x
Common stock=10000
dividends paid=5000
80000=40000+10000+x-5000
X=35000